# PINN-MEP: Continuous Neural Representations for Minimum-Energy Path Discovery in Molecular Systems

### v3

@@ 5 Limitations
While our feedforward neural architecture can model conformational transitions, it could benefit from incorporating domain-specific structural priors to improve training efficiency and convergence speed. Physical force fields introduce challenges through high-energy spikes that can cause training instability and divergence in some runs. Our gradient clipping approach mitigates this issue but doesn’t universally prevent training failures.
The method has been validated primarily on small to medium-sized protein systems, and scaling to larger multi-domain proteins will require additional optimization strategies. Our approach is also constrained to systems with differentiable force fields, but this limitation diminishes as more implementations become available.
Finally, while computationally efficient, our method focuses on identifying minimum-energy paths rather than sampling the full statistical ensemble of transition pathways, representing a trade-off between efficiency and mechanistic detail.
@@ 6 Conclusion
We introduced a novel approach to molecular transition path discovery that leverages physics-informed neural networks and differentiable force fields to enable continuous representation of conformational changes. By reformulating minimum-energy path discovery as a neural optimization problem, we demonstrated the method’s effectiveness on both the small AIB9 peptide and the explicitly solvated BPTI protein, showing its ability to discover physically meaningful transition pathways across different complexity scales.
Our approach achieves remarkable computational efficiency, reducing computational requirements by approximately six orders of magnitude compared to traditional MD simulations for the BPTI system. These advances make previously intractable conformational transition studies accessible on standard computational resources.
While our current implementation shows promising results, several directions for improvement remain. A key limitation is the simple feedforward architecture used to represent the transition paths. Future work should explore neural architectures that better respect the structural organization of protein systems. Ramakrishnan et al. (2025) investigates non-uniform sampling schemes along the path, among other innovative design decisions that might further aid convergence. Integrating pretrained generative models suggested by Raja et al. (2025) could provide better initial transition guesses, potentially further accelerating convergence by starting optimization from more plausible conformational pathways. Extending the method to handle larger multi-domain proteins is also a promising frontier. Future work could also explore methods to estimate free energies and reaction rates from our discovered paths, which would provide valuable kinetic and thermodynamic insights beyond the current geometric focus.
@@ 7 Broader Impacts
Our research democratizes access to molecular simulation by enabling transition path studies on standard hardware. This may accelerate drug discovery and protein engineering while contributing to sustainable computing through drastically reduced energy consumption. We see no direct negative societal impacts, though we acknowledge that accelerated protein engineering capabilities could potentially be misused in the future, and we encourage responsible use with appropriate experimental validation.
@@ A.1.1 From Onsager-Machlup to a geometric line integral
The probability density of a trajectory on is given by:
where is the Onsager-Machlup functional:
@@
with equality when is parallel or antiparallel to . This is precisely the defining characteristic of the minimum-energy path.
@@ A.1.2 Discretization and connection to energy differences
To compute this geometric action numerically, we introduce a progress coordinate and a mapping such that and . The relationship between the progress coordinate and arc length is:
@@
Using Taylor’s theorem and assuming , we can relate energy differences to gradient magnitudes:
where is a remainder of order and is the angle between the gradient and displacement vectors.
For a path following the minimum-energy trajectory, the gradient aligns with the path direction ( as ), giving:
@@ A.1.3 Establishing the energy-based loss function
Setting and summing over all segments:
@@
Since , , and are constants, minimizing is equivalent to minimizing the upper bound on the geometric action. This provides the theoretical justification for our energy-based loss function.
@@ A.1.4 Neural network implementation
For practical optimization, we parametrize the path using a neural network with parameters . We employ stochastic sampling of the progress variable to construct our loss function:
where are uniformly sampled progress coordinates.
@@
By minimizing through gradient-based optimization, the neural network learns to represent low-energy pathways connecting the stable states, effectively discovering the minimum-energy path predicted by the Onsager-Machlup framework.
@@ A.2 Connection between Doob’s Lagrangian and Minimum-Energy Path Optimization
This functional quantifies the cost of controlling a stochastic process to achieve desired endpoint conditions, subject to the system’s energy landscape. Intuitively, it measures the amount of "effort" needed to steer the dynamics from starting state A to target state B, with smaller values indicating more probable transition paths.
@@ A.2.1 Comparison with Doob’s Lagrangian Method
We provide a direct comparison between our PINN-MEP approach and the method of Du et al. in terms of computational efficiency, as evaluated by the number of force field/energy function calls, and the maximum energies observed in the paths, corresponding to the peak of the potential barrier. To provide a rigorous comparison, we conducted benchmarks using experimental conditions identical to those reported by Du et al. for alanine dipeptide: AMBER14 force field, 300K temperature. Table 1 presents these results alongside those reported in Du et al. (2024) and their recent follow-up work lee_scaling_2025. The only exception is that we do not evaluate trajectory probability, as the assumption of optimization over a non-physical path length makes the calculation of this quantity impossible.
@@ Appendix B Comparison with Other Related Methods
B.0.1 Network Architecture
Model type: Transition MLP with sinusoidal time embedding
Network depth: 10 hidden layers
Network width: 8 width factor (scaled relative to the number of atoms)
Time embedding: Sinusoidal position encoding with maximum period of 100,000
Activation function: ReLU
B.0.2 Training Parameters
Optimizer: Adam
Learning rate: 0.0004
Batch size: 1
Frames per batch: 8
Spring constant (): 0.0
Interpolation method: Linear
@@ B.2.3 Hybrid Neural Network Architecture
Protein representation: Neural network with sinusoidal time embedding for the protein atoms (58 residues, 892 atoms)
@@
Control points: 5 control points for the spline interpolation
@@ B.2.4 Training Parameters
Random seed: 1
B.2.5 Force Field Parameters
@@
B.2.6 Computational Performance
For the BPTI system with explicit solvent (approximately 11,000 atoms total):
Training time: Approximately 15 minutes on a single NVIDIA A6000 GPU

### v4

@@ 5 Limitations
While our feedforward neural architecture can model conformational transitions, it could benefit from incorporating domain-specific structural priors to improve training efficiency and convergence speed. Furthermore, physical force fields introduce challenges through high-energy spikes that can cause training instability and divergence in some runs. Our gradient clipping approach mitigates this issue but doesn’t universally prevent training failures.
The method has been validated primarily on small to medium-sized protein systems, and scaling to larger multi-domain proteins could require additional optimization strategies. Our approach is also constrained to systems with differentiable force fields, but this limitation diminishes as more implementations become available.
Finally, while computationally efficient, our method focuses on identifying minimum-energy paths rather than sampling the full statistical ensemble of transition pathways, representing a trade-off between efficiency and exhaustively sampling a system.
@@ 6 Conclusion
We introduced a novel approach to molecular transition path discovery that leverages physics-informed neural networks and differentiable force fields to enable the continuous representation of conformational changes. By reformulating minimum-energy path discovery as a neural optimization problem, we demonstrated the method’s effectiveness on both the small AIB9 peptide and the explicitly solvated BPTI protein, showing its ability to discover physically meaningful transition pathways across different complexity scales.
Our approach achieves remarkable computational efficiency, reducing computational requirements by approximately six orders of magnitude compared to traditional MD simulations for the BPTI system. While our method yields a single pathway rather than the complete ensemble of transition trajectories, this represents a valuable trade-off that makes previously intractable conformational transition studies accessible on standard computational resources.
While our current implementation shows promising results, several directions for improvement remain. Future work should explore neural architectures that better respect the structural organization of protein systems. Ramakrishnan et al. (2025) investigates non-uniform sampling schemes along the path, among other innovative design decisions that might further aid convergence. Integrating pretrained generative models suggested by Raja et al. (2025) could provide better initial transition guesses, potentially accelerating convergence by starting optimization from more plausible conformational pathways. Extending the method to handle larger multi-domain proteins is also a promising frontier. Future work could also explore methods to estimate free energies and reaction rates from our discovered paths.
@@ 7 Broader Impacts
Our research democratizes access to the study of larger molecular systems by enabling transition path studies on standard hardware. This may accelerate drug discovery and protein engineering while contributing to sustainable computing through drastically reduced energy consumption. We see no direct negative societal impacts, though we acknowledge that accelerated protein engineering capabilities could be misused in the future, and we encourage responsible use with appropriate experimental validation.
@@ A.1 Extended derivation of the energy-based loss function
In this appendix, we provide a comprehensive derivation of our energy-based loss function for discovering minimum-energy paths (MEPs) between molecular configurations. The derivation proceeds in four main steps: (1) establishing the path probability density from overdamped Langevin dynamics, (2) reformulating this probability in terms of the Onsager-Machlup action functional, (3) simplifying this functional to a geometric line integral, and (4) developing a discretized approximation suitable for neural network optimization.
Our goal is to find the most probable transition path between two stable molecular configurations and . This path corresponds to the minimum-energy pathway through the potential energy landscape and represents the most likely mechanism of conformational change.
Assumptions and Approximations
The derivation relies on several key assumptions:
The molecular system evolves according to overdamped Langevin dynamics, appropriate for high-friction biomolecular environments
The path endpoints and are fixed and correspond to stable states
The potential energy function is differentiable everywhere along the path
We can represent the path as a continuous, parameterized curve in configuration space with sufficient flexibility
The path’s parameterization allows for effective length changes, permitting the path to adjust its stretching as needed to follow energy gradients optimally
We also make several important approximations during the derivation:
We use a first-order Taylor expansion to relate energy differences to gradient magnitudes
We introduce stochastic sampling to estimate the continuous energy integral
For a minimum-energy path, the displacement tends to align either parallel or anti-parallel with the gradient direction (i.e., for path segments)
We use the triangle inequality to establish an upper bound on the sum of absolute energy differences
While the last two approximations do not change the location of the global minimum, they do affect the optimization landscape before that minimum is reached.
@@ A.1.1 From Onsager-Machlup to a geometric line integral
To derive the path probability, we discretize the time interval into steps of duration , approximating the dynamics as:
where are independent standard Gaussian random variables with zero mean and unit variance.
The conditional probability of transitioning from to follows a Gaussian distribution:
where is the dimensionality of the system. The probability of the entire discretized path is given by the product of these conditional probabilities:
Taking the limit as and with fixed, this sum approaches a path integral:
Thus, the path probability can be written as:
The exponent term is recognized as the Onsager-Machlup action functional:
This action functional penalizes paths that deviate from following the force . Maximizing the probability of a transition path (subject to fixed endpoints and ) is equivalent to minimizing this action functional.
@@
with equality when is parallel or anti-parallel to , with the direction determined by whether the path is ascending or descending the energy landscape. This is precisely the defining characteristic of the minimum-energy path (MEP) - a path that follows the potential energy landscape’s gradient while avoiding high-energy regions.
@@ A.1.2 Discretization and connection to energy differences
To compute this geometric action numerically, we introduce a progress coordinate and a mapping such that and , where represents the network parameters. The relationship between the progress coordinate and arc length is:
@@
Using Taylor’s theorem, we can relate energy differences to gradient magnitudes:
where is the remainder of the order and is the angle between the gradient and displacement vectors.
For a path following the minimum-energy trajectory, the gradient aligns with or against the path direction ( as , depending on whether the path is ascending or descending the energy landscape), giving at the minimum of the optimization:
@@
This forms a surrogate objective with the same minimum as our original action.
@@ A.1.3 Establishing the energy-based loss function
Choosing to be the minimum energy along the entire path and summing over all segments:
@@
Since , , and are constants during optimization, minimizing is equivalent to minimizing the upper bound on the geometric action. This provides the theoretical justification for our energy-based loss function.
@@ A.1.4 Neural network implementation
For practical optimization, we parametrize the path using a neural network with parameters . To implement our approach, we sample points along the path and minimize the energy at these sampled locations. This represents a transition from the theoretical derivation using discrete segments to a practical implementation using sampled points for more efficient optimization. We employ stochastic sampling of the progress variable to construct our loss function:
where are uniformly sampled progress coordinates and is the batch size (the number of points sampled along the path during each optimization step).
@@
By minimizing through gradient-based optimization, the neural network learns to represent low-energy pathways connecting the stable states, effectively discovering the minimum-energy path predicted by the Onsager-Machlup framework. The resulting path follows the potential energy landscape’s gradient while avoiding high-energy regions, representing the most probable transition mechanism in the overdamped limit.
@@ A.2 Connection between Doob’s Lagrangian and Minimum-Energy Path Optimization
This functional quantifies the cost of controlling a stochastic process to achieve desired endpoint conditions, subject to the system’s energy landscape. Intuitively, it measures the amount of "effort"/action needed to steer the dynamics from starting state A to target state B, with smaller values indicating more probable transition paths.
@@ A.2.1 Comparison with Doob’s Lagrangian Method
We provide a direct comparison between our PINN-MEP approach and the method of Du et al. in terms of computational efficiency, as evaluated by the number of force-field/energy function calls, and the maximum energies observed in the paths, corresponding to the peak of the potential barrier. To provide a rigorous comparison, we conducted benchmarks using experimental conditions identical to those reported by Du et al. for alanine dipeptide: AMBER14 force field, 300K temperature. Table 1 presents these results alongside those reported in Du et al. (2024) and their recent follow-up work Lee et al. (2025). The only exception is that we do not evaluate trajectory probability, as the assumption of optimization over a non-physical path length makes the calculation of this quantity impossible.
@@ B.1.3 Network Architecture
B.1.4 AIB9 Network Architecture
Model type: Transition MLP with sinusoidal time embedding
Network depth: 6 hidden layers
Network width: 4 width factor (scaled relative to the number of atoms)
Time embedding: Sinusoidal position encoding with maximum period of 1,000,000
Activation function: ReLU
B.1.5 Training Parameters
Optimizer: Adam
Learning rate: 0.00001
Batch size: 4
Frames per batch: 32
Spring constant (): 0.0
Interpolation method: Linear
@@ B.2.3 Hybrid Neural Network Architecture
Protein representation: Neural network with 6 layers, a width factor of 4, and sinusoidal time embedding for the protein atoms (58 residues, 892 atoms)
@@
Control points: 7 control points for the spline interpolation
@@ B.2.4 Training Parameters
B.2.5 Handling High-Energy Spikes During Training
Physical force fields present unique optimization challenges due to occasional extreme energy spikes when atoms approach one another too closely. To address this, we implemented adaptive gradient clipping in our training procedure:
where is a clipping threshold (set to in our experiments). We specifically use adaptive gradient clipping rather than a fixed threshold because our hybrid architecture contains two fundamentally different parameter types: neural network weights for protein coordinates and spline control points for water molecules. The adaptive nature of the clipping automatically adjusts to the appropriate scale for each parameter type.
Additionally, we apply different scaling factors to these parameter groups:
with for water-related parameters. This differential treatment further balances the influence of each component during optimization. While this approach does not completely eliminate training instabilities, we find it sufficient for most protein systems when combined with careful initialization and learning rate scheduling.
B.2.6 Force Field Parameters
@@
B.2.7 Computational Performance
For the BPTI system with explicit solvent (approximately 8,300 atoms total):
Training time: Approximately 7 to 15 minutes on a single NVIDIA A6000 GPU